# Ranking Unraveled: Recipes for LLM Rankings in Head-to-Head AI Combat

### v1

@@ 2.4 Glicko Rating system
2.5 Markov Chain
The Markov Chain Model is a nonparametric ranking algorithm adapted from Callaghan et al. (2003, 2004) and adapted into a Markov Chain Model by Kvam and Sokol (2006). The Markov chain employs a series of random walkers traversing a graph, with each node representing a player/model in the tournament connected by an edge denoting a match between two nodes. Each walker chooses an initial node $i$, staying in the node or moving to an adjacent node $j$ according to a single condition $p$. In a random match between $i$ and $j$, the walker moves to the winning node with a probability of $p$ for $p>0.5$, a hyperparameter chosen in advance. Every time a walker moves or stays on a node, it represents a vote for that node, and the sum of all votes from every walker represents the rank of that player/model. In practice, the optimal value for $p$ can vary depending on the tournament. In our experiments, we use $p=0.8$ as the default value similar to Callaghan et al. (2003), which achieved stable results with $p$ ranging from $p=0.75$ to $p=0.95$. When expressed as a Markov chain, the elements $T_{ij}$ of the transition matrix $T$ are the probabilities that a walker moves from model $i$ to $j$ in a single time step, with the initial values computed as follows:
@@ 2.5 Markov Chain
3 Desirable Properties of Ranking System
The fundamental goal of a ranking system is to provide a structured way to compare and order items based on their performance or quality. In the context of LLMs, an effective ranking system allows researchers, practitioners, and end-users to identify the most capable models for their specific needs. It simplifies decision making and optimizes resource allocation by highlighting models that are likely to deliver superior performance in real-world applications. However, employing such involves addressing several key properties to ensure its validity and reliability. In this section, we identify and elaborate on three essential properties that a ranking system for LLMs should adhere to: 1) Transitivity, 2) Prediction Accuracy, and 3) Sensitivity to Hyperparameters and Battle Conditions.
3.1 Transitivity
Transitivity is a fundamental property for any ranking system to maintain consistency. Transitivity is defined as follows:
Definition 3.1.
For any three models $i,j,k\in M$; their match-ups are transitive if $p_{ij}>0.5,p_{jk}>0.5$ and $p_{ik}>0.5$. If any three models $i,j,k$ are transitive by definition 3.1, we say that a ranking system preserves transitivity if it ranks these models in order $i,j,k$. Transitivity ensures that the ranking is coherent and interpretable.
In the context of LLMs, transitivity guarantees that the relative strengths of models are accurately reflected across various pairwise comparisons. When utilized by researchers and practitioners, the general assumption of a model’s placement on a leaderboard such as the LMSYS Chatbot Arena Leaderboard Chiang et al. (2024) indicates its overall strength compared to its counterparts. This means that a lack of transitivity can lead to paradoxical rankings, where the ordering of models becomes circular or inconsistent, thereby diminishing the trust in the ranking system.
3.2 Prediction Accuracy
Like transitivity, prediction accuracy is a critical metric that assesses how well the ranking system forecasts the results in head-to-head matches. It measures the probability that the ranking system correctly predicts the winner between any two LLMs. High prediction accuracy indicates that the ranking system is reliable and can be trusted to make informed decisions about the superiority of the model for a downstream task. For combat-style evaluation in LLM, where assessment is based on subjective human judgment, prediction accuracy acts as an indicator of the ranking system’s ability to align with human preferences. This alignment is crucial for tasks like open-ended text generation or conversational agents, where the model’s performance is best judged by human preference. Thus, the ranking system must reflect statistical superiority and resonate with qualitative human evaluations to truly enhance the utility and adoption of LLM.
3.3 Sensitivity to Hyperparameters and Battle Conditions
4 Evaluation
Algorithm Arena (%) SLAM (%)
Elo 68.24 52.5
Markov 51.38 51.67
Glicko 56.54 53.33
Bradley-Terry 77.29 56.67
4.1 Datasets
To evaluate these ranking systems, we use two diverse human evaluation datasets representative of two evaluation styles (Arena Style vs Controlled Style):
These two diverse datasets allow us to evaluate the ranking algorithms under different conditions, providing a comprehensive perspective on their behavior.
4.2 Measuring Transitivity on Human Feedback Data
Using the Chatbot Arena and SLAM datasets, we evaluated the ability of each ranking algorithm to preserve transitivity. Recall that transitivity as defined in 3.1 is the property that if $i$ is preferred to $j$ and $j$ is preferred to $k$, then $i$ should be preferred to $k$. We evaluated the transitive performance of each ranking algorithm by first calculating the number of triples within the data set. Three models $i,j,k$ are part of a triple if $i$ wins the majority of comparisons with $j$, and $j$ wins the majority of comparisons with $k$, and transitivity is considered preserved if this order is maintained in the final ranking. Table 1 shows the performance of each ranking algorithm under the arena and controlled style evaluation scenarios. We observe that the Bradley-Terry model is the most performant, preserving transitivity in $77.29\%$ and $56.67\%$ of the triples, respectively, compared to its closest competitor, Elo, at $68.24\%$ and $52.5\%$.
Algorithm Arena (F1) SLAM (F1)
Elo 0.87 0.90
Markov 0.88 0.77
Glicko 0.88 0.88
Bradley-Terry 0.87 0.82
Table 2: Ranking Algorithm Performance (F1 score) on Chatbot Arena & SLAM Dataset.
4.3 Prediction Accuracy
4.3.1 Performance on the Arena Dataset
Due to its uneven distribution of model match-ups, the Arena data set presents an interesting challenge for ranking algorithms to predict future outcomes. When matches between models are unevenly distributed, some classification algorithms struggle with prediction accuracy. In Figure 4, our results show the Elo model as the most accurate, achieving an F1 score of $0.90$ compared to the F1 score of $0.88$, $0.82$, and $0.77$ for the models Glicko, Bradley-Terry, and Markov Chain, respectively (Table 2).
The impact of the match-up distribution on the prediction accuracy is shown in the performance of the Markov chain which is affected by the sparsity of certain model match-ups, leading to lower prediction accuracy as shown by the poor performance on models such as deepseek-llm-67b-chat and dolphin-2.2.1-mistral-7b in figure 4 of the appendix.
Another example of the impact of matchup distribution is shown in the performance of the Bradley-Terry model in appendix figure 4. We observe that the Bradley-Terry system suffers when dealing with "powerful models," where there is a strong imbalance between a model’s wins and losses. For example, in the arena dataset, gpt-4-turbo had a win/loss ratio of 12288/3979. This skew in model wins/loss results in overestimation of the model’s strength as shown by the low F1 score of $0.82$ achieved by gpt-4-turbo using Bradley-Terry compared to an F1 score of $0.95$, $0.96$ and $0.92$ from Elo, Glicko, and Markov Chain, respectively. This is known as the ‘rare events’ problem in logistic regression King and Zeng (2001). One method of countering this is to use a weighted logistic regressionChiang et al. (2024). We observe however, that this has a negligible impact on the final result.
4.3.2 Prediction accuracy on the SLAM dataset
4.4 Hyperparameter Sensitivity
To measure the sensitivity of the hyperparameters, we calculate the F1 score of each ranking system as described in Section 4.3, varying each hyperparameter over a range of 100 values. We then plot the distribution of F1 scores for each ranking algorithm across the two datasets. Figure 2 showcases the range of results produced by Elo, Markov and Glicko across the two evaluation scenarios.
Method Elo Markov Glicko Bradley-Terry Win-Rate
Elo Rating 1.00/1.00 0.74/0.89 0.86/0.93 0.94/0.95 0.93/0.91
Markov - 1.00/1.00 0.76/0.99 0.81/0.97 0.78/0.98
Glicko - - 1.00/1.00 0.81/0.99 0.89/0.99
Bradley-Terry - - - 1.00/1.00 0.91/0.98
Win-Rate - - - - 1.00/1.00
Figure 2: The distribution of F1 scores for SLAM (left) and ARENA (right) datasets. Each line represents the distribution of F1 scores for 100 different hyper-parameters for the Elo, Markov, and Glicko algorithms across a subset of models.
4.4.1 Elo k-factor
The $k$ factor in Elo dictates the extent to which the score of a player is influenced by the outcome of an individual match. For example, with large $k$ values, highly ranked players can incur a significant penalty when losing to a player with a lower ranking, while defeating such a player only results in a slight increase in score. In sports games, the strength of players fluctuates over time, influenced by their experience and age; on the other hand, in llm evaluation the models tend to remain unchanged during their lifespan. [^3] Models that undergo retraining or fine-tuning are usually considered new models The evaluation of the performance of the model (depicted in Figure 2) at varying levels of hyperparameter $k$ shows marked sensitivity, which emphasizes the need for meticulous tuning. Elo shows the largest variation across both datasets, and is less stable, particularly for the smaller SLAM dataset. Through our tuning of this parameter, we determined that the ideal $K$ factor for Elo is less than the standard 32 commonly used in previous research Dettmers et al. (2024); Wu and Aji (2023); Chiang et al. (2023). This reduced convergence rate can be attributed to the complexities of human preferences as opposed to the typical assessment of real-world games, which rely on the differing abilities of opposing teams.
4.4.2 Markov Chain: $p$ Score
When using Markov chain, in a random match between $i$ and $j$, the walker moves to the winning node with a probability of $p$ for $p>0.5$, a hyperparameter chosen in advance. Hence, if we set a value for $p$ closer to 0.5, then models with lower win rates may outrank those with higher win rates. However, for the Arena data set, choosing a higher $p$ value will cause the algorithm to prioritize the win rate over the number of games played; for example, a model with 1000 games with a high win rate may be ranked higher than a model with 30,000 games and a slightly lower win rate. We observe that Markov shows stable performance on the controlled SLAM datasets, performing as well as Glicko (Figure 2), however, the algorithm struggles on the large ARENA dataset due to its over reliance on win-rates for calculating ranks. When optimizing the value of this parameter, we find that the optimal $p$ value for the Markov chain on the SLAM dataset is larger than the Arena dataset, indicating that the model’s strength is greatly influenced by its overall win rate.
4.4.3 Glicko: Rating deviation
For Glicko, the ranking deviation $(\sigma)$ is a measure of consistency or variability in a player’s performance over time and determines how flexible a player’s rating can be. The volatility itself is updated as part of the rating calculation, making it dynamic and player-specific. However, the initial value and the formula used to adjust it are defined by the system and can be modified based on the specific application. If volatility is initially set too high, the system will interpret early game results as more variable or uncertain. This leads to larger rating adjustments, making the player’s rating fluctuate more. This can cause new or improving players to initially overreact to wins or losses, potentially skewing their rating and making it difficult to stabilize. In contrast, setting the volatility too low initially restricts the system from adjusting the rating freely. As a result, players, especially newcomers, may take much longer to reach a rating that accurately reflects their true skill level. This is particularly impactful in fast-changing or skill-evolving games, where a low initial volatility setting could lead to underestimation of skill. This high volatility further affects the ability of the algorithm in prediction accuracy of the algorithm. We observe that Glicko’s prediction performance is consistent across both datasets, showing that its performance is less impacted by changes in its hyperparameters compared to Elo and Markov Chain. We theorize that this is due to the large number of games played for each model in both scenarios, allowing the system to adjust the rating deviation dynamically.
5 Discussion
Criteria Elo Markov Glicko Bradley-Terry
In the sections 3 and 4, we outlined the desired properties of ranking systems and presented experiments assessing the performance of these systems in achieving these criteria. Based on these observations, this section puts these results into perspective and provides a set of practical recommendations for selecting the most appropriate ranking algorithm. We outline which algorithms perform best under specific circumstances, offering guidance on their optimal use depending on the structure of the dataset, the matchup distribution, and the model evaluation goals. These recommendations, summarized in Table 4, aim to improve the reliability and applicability of classification systems in real-world LLM evaluations.
5.1 Challenges with Elo
An important result in this paper is that Elo, even with $>$ 1,000 permutations, struggles to achieve stable rankings. This indicates that Elo’s dependency on permutations for ranking stabilization is inadequate, challenging previous studies Boubdir et al. (2023). Combined with the high sensitivity to hyperparameters and the average performance in preserving transitivity. We do not recommend ELO as an algorithm to rank LLM performance.
5.2 Choose Bradley-Terry for Small Controlled Datasets
In section 4.3, we demonstrate that ranking algorithms exhibit varied performance based on battle conditions. When matches are evenly distributed, ranking algorithms produce similar results. Whereas, a heavily skewed distribution can negatively impact some methods such as Elo and Bradley-Terry. In this scenario where matchups are balanced and the number of matches are small, objectiveur findings show that Bradley-Terry is best due to its ability to preserve transitivity while also maintaining comparable prediction accuracy.
5.3 Choose Glicko for large uneven datasets
For larger datasets, especially those with uneven matchup distributions, the Bradley-Terry model encounters difficulties due to the rare event problem. This problem occurs when new models with high win rates, participate in fewer matchups than well-established models. As a result, these new models can be ranked disproportionately higher based on limited data, leading to biased rankings. In this scenario, the Glicko system provides a superior solution. Glicko utilizes a rating deviation parameter to account for uncertainty in a model’s ranking due to limited matchups. This allows the system to adjust the rankings more conservatively for models with fewer comparisons, preventing them from being ranked too high based only on a small number of matches. As data sets become larger and more uneven, Glicko’s ability to incorporate rating deviation ensures that the rankings accurately reflect true performance throughout the dataset, even when new models appear with limited match histories. This flexibility makes Glicko particularly effective for large-scale LLM assessments where uneven matchup distributions are prevalent.
5.4 Applicability of Bradley-Terry for other Dataset Types
While Bradley-Terry model is particularly well suited for small, evenly distributed datasets, there are compelling reasons to use it in other scenarios, such as small uneven datasets or large, evenly distributed datasets. For small uneven datasets, Bradley-Terry’s simplicity and ability to model pairwise comparisons make it a robust choice, even when the data distribution is skewed. It does not require parameter estimation, which allows it to handle small datasets efficiently without the need for complex adjustments or hyperparameter tuning, making it ideal for situations where uneven distributions might otherwise require additional preprocessing.
In the case of large, evenly distributed datasets, Bradley-Terry can still be effective because of its scalability and ease of interpretation. For large data sets where the distribution of results is balanced, Bradley-Terry’s ability to produce clear, interpretable rankings without adding excessive computational overhead is a significant advantage. Although more complex models such as Glicko can offer finer granularity in cases of large uneven data, Bradley-Terry’s efficiency and interpretability make it a viable alternative when computational simplicity and transparency are prioritized.
6 Related Work
The evaluation of LLMs is a multifaceted process that involves two key components: the tasks assigned to the model and the strategy used to measure the model’s performance on those tasks. The choice of tasks and evaluation strategies can significantly impact the reliability and validity of the evaluation results. In this paper, we focus specifically on the ranking of LLMs based on pairwise evaluations and evaluate the performance of traditional ranking algorithms in this context.
Pairwise ranking has recently emerged as a popular method for evaluating LLMs, as it allows for direct comparisons between models and provides a more nuanced assessment of their performance than traditional metrics Chiang et al. (2024). This evaluation involves presenting human judges with pairs of model outputs and asking them to select the better one. The outcomes of these pairwise comparisons are then aggregated into final rankings using ranking algorithms such as Elo Elo (1978). Despite their widespread use, there has been limited research on the applicability of these ranking algorithms to LLM evaluation, especially in scenarios where models are static and datasets are unevenly distributed. Works such as Boubdir et al. (2023) have critically analyzed the compatibility of Elo, the most popular and widely used algorithm adapted for this use case, with LLM evaluations by examining how well it adheres to reliability and transitivity. However, Elo is not the only ranking algorithm used in LLM evaluation. There exists a plethora of other algorithms that have been proposed for ranking in different contexts, such as the Bradley-Terry model Bradley and Terry (1952), which is widely used in sports rankings, and the Glicko model Glickman and Jones (1999), which is used in chess rankings. These algorithms have been shown to outperform Elo in certain scenarios, but their performance in the context of LLM evaluation has not been systematically evaluated. In addition, there are several factor that can affect the performance of these algorithms, such as the distribution of the data, the number of models being compared, and the number of pairwise comparisons. As such there is a need for a systematic comparison of ranking algorithms in the context of LLM evaluation to identify the most suitable algorithm for different scenarios which this paper aims to address.
7 Conclusion
8 Limitations
While our study offers a comprehensive exploration into the ranking of Large Language Models (LLMs) through head-to-head combat, several limitations warrant discussion.
Scalability Constraints
One significant limitation arises from the scalability constraints inherent in pairwise evaluations. As the number of LLMs in the ecosystem continues to expand, the number of required comparisons grows quadratically. This introduces computational challenges and resource constraints that may limit the feasibility of exhaustive head-to-head evaluations, particularly for larger sets of models.
Human Feedback Variability
Another concern is the variability in human feedback, which our rankings are heavily based on. Human judgments can be subjective and influenced by numerous factors, including the individual’s background, expertise, and contextual understanding. This variability can introduce noise into the ranking system, which can affect the stability and reliability of the evaluations.
Appendix A Appendix
A.1 F1 scores per model

### v2

@@ A.1 Glicko Algorithm Details
A.2 Markov Chain Details
The Markov Chain Model was originally developed as a random walker model Callaghan et al. (2003, 2004) where ranks were calculated by solving a series of differential equals. This method was adapted into a Markov Chain Model by Kvam and Sokol (2006). When expressed as a Markov chain, the elements $T_{ij}$ of the transition matrix $T$ are the probabilities that a walker moves from model $i$ to $j$ in a single time step, with the initial values computed as follows:
@@ A.2 Markov Chain Details
A.3 F1 scores per model
@@ A.4.2 Procedure
A.5 Elo’s Instability at higher permutation
To analyze the performance of Elo at higher permutations, we conducted two separate experiments on the SLAM dataset.
A.5.1 Experimental Setup
For a single $K$ factor of 25, we evaluated Elo rankings by performing two separate evaluations of each $10,000$ permutation using the strategy of the previous section.
A.5.2 Procedure
We observed the ranks produced by each experiment, showing that even with a large permutation, ELO can produce two different ranks for the same two models6.
Figure 6: Two separate ranks produced by ELO for permutations of 10,000 each.
A.6 Evaluating the consistency of crowd-workers evaluators
In this section, we recommend a method to evaluate the consistency of crowd-worker evaluations based on the Analytic hierarchy process (AHP) practices. We begin by formally defining cycles in a pairwise ranking dataset.
Definition A.1.
For models $i_{1},i_{2},\cdots i_{k}\in M$, if $p_{i_{1}i_{2}}>0.5,p_{i_{2}i_{3}}>0.5,\cdots,p_{i_{k}i_{1}}>0.5$, then their games is said to be cyclic. A cycle of length three is called a Rock, Paper, Scissors (RPS) cycle.
For any three models $i,j,$ and $k$, it is expected that if a reviewer prefers $i$ over $j$ and $j$ over $k$, they prefer $i$ over $k$ (transitivity). However, this may not always be due to personal preference, evaluation error, or evaluator negligence. If the transitive property does not hold, we say that the three models are intransitive and the evaluation is inconsistent.
Measuring the consistency of pairwise ranking falls under the domain of AHP. According to AHP theory, if the number of RPS cycles remains small, it reflects the complexity of the ranking problem rather than the judges’ complacency. However, a high number of cycles represents a low quality of judgments.
According to Kendall and Smith (1940), the expected number of RPS cycles for a pairwise comparison dataset of $m$ items follow a $\chi^{2}$ distribution, and if the actual number of RPS cycles is significantly less than the expected number, we can be reasonably confident that the dataset is consistent.
The number of RPS cycles can be calculated by first generating the win rate matrix $P$ as defined in section 4.2. This is followed by the adjacency matrix $A$, where $a_{ij}=1$ if $p_{ij}>0.5$ and 0 otherwise. Finally, the total number of RPS cycles is calculated as $\rho=1/3\sum_{i=1}^{M}a^{*}_{i,i}$, where $a^{*}_{ii}$ are diagonals of $A^{3}$anjaya et al. (2022).
The $\chi^{2}$ statistic is calculated as:
$\chi^{2}=\frac{8}{k-4}\left\{\frac{1}{4}\binom{k}{3}-d-\frac{1}{3}+\rho\right\}$ (11)
$v=\frac{k(k-1)(k-2)}{(k-4)^{2}}$ (12)
where $v$ is the degrees of freedom.
The RPS count and $\chi^{2}$ statistic for both the SLAM and Arena datasets are shown in table 5. At a p-value close to 0, we can be confident that inconsistencies are not due to the quality of the judges.
Table 5: Non-Transitivity statistic for the two datasets that were used. The p-value is calculated based on the $\chi^{2}$ estimated via Kendal’s Chi-squared approximation to the circular triad distribution.
Dataset RPS Count $\chi^{2}$ p-value
SLAM 19 45.06 0.001
Arena 393 1107.25 0.000